OpenStudy (eric_d):

Find coordinates of stationary pt.of inflexion on curve given by eqn y=3x^4- 4x^3

OpenStudy (eric_d):

dy/dx = 12x^3 - 12x^2

OpenStudy (eric_d):

dy/dx=0 12x^3-12^2 12x^(x-1)=0 x=0,1 When x=1, y=-1 Wwhen x=0, y=0 From dy/dx, the points are (0,0) and (1,-1)

OpenStudy (eric_d):

d^2y/dx^2= 36x^2 - 24x 36x^2 - 24x = 0 12x(3x-2)=0 x=0,2/3 When x=0, y=0 When x=2/3 , y= -16/27 When x=0, d^2y/dx^2=36(0)^2-24(0) =0 When x=2/3 , d^2y/dx^2 = 36(2/3)^2 - 24(2/3) = 0 The pt.of inflexion are (0,0) and (2/3 , -27/16)

OpenStudy (eric_d):

Kinda confuse with my own working.. !

OpenStudy (eric_d):

@ikram002p @ganeshie8 @dan815

OpenStudy (eric_d):

So, my working is correct ? @ganeshie8

OpenStudy (eric_d):

@ganeshie8

ganeshie8 (ganeshie8):

so we have 1) dy/dx = 0 at the points `(0,0) and (1,-1)` 2) d^2y/dx^2 = 0 at the points `(0,0) and (2/3 , -27/16)`

ganeshie8 (ganeshie8):

lets find out stationary and inflection points now

OpenStudy (eric_d):

Okay

OpenStudy (eric_d):

How ? Sub value of x into dy/dx and d^2y/dx^2 respectively ?

ganeshie8 (ganeshie8):

First of all what is a stationary point ?

OpenStudy (eric_d):

It's a point whereyby the gradient of curve is always 0

OpenStudy (eric_d):

whereby*

ganeshie8 (ganeshie8):

http://mathworld.wolfram.com/StationaryPoint.html

ganeshie8 (ganeshie8):

Right, so our stationary points are simply the points at which first derivative is 0 : ` (0,0) and (1,-1)`

OpenStudy (eric_d):

Okay

ganeshie8 (ganeshie8):

lets find out inflection points

OpenStudy (eric_d):

Kk

OpenStudy (eric_d):

By using d^2y/dx^=0 rite ?

OpenStudy (eric_d):

dx^2 *

ganeshie8 (ganeshie8):

yes second derivative is 0 for inflection points but not all points where second derivative is 0 is an inflection point

ganeshie8 (ganeshie8):

its like saying : all dogs have four legs but not all four legged pets are dogs

ganeshie8 (ganeshie8):

we have d^2y/dx^2 = 0 at the points `(0,0) and (2/3 , -27/16)` that just tells us that inflection occurs only at `few or all` of these points

ganeshie8 (ganeshie8):

how do we check whether they are inflection points ?

OpenStudy (eric_d):

By testing the sign of d^2y/dx^2 .. : Sub value of x into it

ganeshie8 (ganeshie8):

BINGO!

ganeshie8 (ganeshie8):

sub a value of x to the left and right of the x value of possible point

ganeshie8 (ganeshie8):

to check whether the point `(0,0)` is really an inflection point or not : 1) plugin x = -0.1 into d^2y/dx^2 2) plugin x = +0.1 into d^2y/dx^2 both should give you opposite sign numbers

OpenStudy (eric_d):

So, that applies to point (2/3, -16/27)

OpenStudy (eric_d):

One is positive the other is negative for both the points.

OpenStudy (eric_d):

ince, there's a change of sign, it's a pt of inflexion

OpenStudy (eric_d):

Since *

ganeshie8 (ganeshie8):

Excellent! that means both are inflection points!

OpenStudy (eric_d):

And point (1,1) is a min point coz dy/dx=12 >0 when x=1

OpenStudy (eric_d):

Thank you @ganeshie8

ganeshie8 (ganeshie8):

Yep! and (0,0) is an inflection point so its neither minimum nor maximum

OpenStudy (eric_d):

Okay :)